What is the answer to this equation: 0.6x-7=23?
Answer by Theo(13342)

.6x - 7 = 23

add 7 to both side of the equation to get .6x - 7 + 7 = 23 + 7

combine like terms to get .6x = 30

divide both sides of the equation by .6 to get .6x / .6 = 30 / .6

simplify to get x = 50

to test this out, replace x with 50 in the original equation.

you will get .6 * 50 - 7 = 23

simplify to get 30 - 7 = 23

simplify further to get 23 = 23

this confirms the solution is correct.